A rectangle has a length 2 meters less than twice its width. When 5 meters is added to the width, the result is a square with an area of 144 square meters. Find the dimensions of the original rectangle
Answer by checkley77(12844)

L=2W-2
2W-2=W+5
2W-W=7
W=7 ANS. FOR THE ORIGINAL WIDTH
W+5=12 ANS. FOR THE SIDES OF THE SQUARE.
L=2*7-2
L=14-2
L=12
PROOF:
12*12=144